OpenStudy (anonymous):

add 7m/16 + (-m/16), reduce if possible. a)6m/16 b)3m/8 c)7/16 d)-m/2 *please just say a,b,c,d

OpenStudy (anonymous):

\[\frac{ 7m }{ 16 } + (-\frac{ m }{ 16 }) = \frac{ 7m }{ 16 } - \frac{ m }{ 16 } = \frac{ 7m - m }{ 16 } = \frac{ 6m }{ 16 } \] now you just have to divide both numerator (upper number in fraction) and denominator (lower number in fraction) with 2 and you have your answer;) *sorry, but rules forbid us of giving out complete answers. we have to animate you to participate in process of solving the assignment

OpenStudy (anonymous):

thank you!!